# NashOpt - A Python Library for Computing Generalized Nash Equilibria

### 07 Mar 2026 v2

@@ Abstract
NashOpt is an open-source Python library for computing and designing generalized Nash equilibria (GNEs) in noncooperative games with shared constraints and real-valued decision variables. The library exploits the joint Karush-Kuhn-Tucker (KKT) conditions of all players to handle both general nonlinear GNEs and linear–quadratic games, including their variational versions. Nonlinear games are solved via nonlinear least-squares formulations, relying on JAX for automatic differentiation. Linear–quadratic GNEs are reformulated as mixed-integer linear programs, enabling efficient computation of multiple equilibria. The framework also supports inverse-game and Stackelberg game-design problems. The capabilities of NashOpt are demonstrated through several examples, including noncooperative game-theoretic control problems of linear quadratic regulation and model predictive control. The library is available at https://github.com/bemporad/nashopt.
@@ 1 Introduction
Generalized games describe situations where several agents take decisions by optimizing individual objective functions under common constraints on their decision variables. Examples include traffic participants using the same roads, wireless systems sharing limited bandwidth, energy systems where prosumers share transmission capacity and balance constraints [10, 38], or firms competing in a market [3]. The solution concept for these problems, in which the cost function and the feasible set of each agent depends on the decisions taken by the other agents, is the generalized Nash equilibrium (GNE), a state in which no agent has an incentive or the possibility to change their decision [2].
GNEPs have been widely studied in the literature, with various algorithms proposed for their solution, including splitting-based operator frameworks [39, 3], approaches built on augmented Lagrangian formulations [34, 27], interior-point optimization strategies [12], Newton-type algorithms [16], distributed algorithms [39, 37, 3], and learning-based methods only relying on measuring best responses [14]. For the specific class of GNEPs in which the shared constraints are linear, efficient solution methods exist for linear objectives [13, 20], multiparametric quadratic programming for strictly convex quadratic objectives [24], and distributed payoff-based algorithms for convex games [36].
@@ 1.1 Contribution
Motivated by the relative scarcity of software packages for solving rather general classes of GNEPs and their application to game-theoretic control problems, in this paper we describe the mathematical formulations of our package NashOpt, a Python library for solving GNEPs based on satisfying the KKT conditions of each player jointly. The library supports nonlinear GNEPs, which are solved via nonlinear least-squares methods, and linear-quadratic GNEPs, which can be solved more efficiently to global optimality via mixed-integer linear programming (MILP) and can be used to enumerate multiple equilibria for the same game corresponding to different combinations of active constraints at optimality. We also address game-design problems, where a central authority can optimize some parameters of the game to achieve a desired equilibrium in accordance with the Single-Leader-Multi-Follower (SLMF) Stackelberg game setting [1]. Moreover, we leverage the GNE solution methods to solve game-theoretic control problems, including non-cooperative linear quadratic regulation (LQR) and model predictive control (MPC) problems [23, 29]. The main purpose of the tool is to analyze and predict the effect of different agents pursuing their own objectives in a shared environment, and to design the game to promote desired equilibrium behaviors.
@@
The library is open-source and available at https://github.com/bemporad/nashopt, and can be easily installed via pip install nashopt.
@@ 2 Problem Formulation
where $x_{i}\in\mathbb{R}^{n_{i}}$ collects the decision variables optimized by agent $i$, $i=1,\ldots,N$, $x_{-i}$ collects those of all other agents, $x=\mathop{\rm col}\nolimits(x_{1},\ldots,x_{N})\in\mathbb{R}^{n}$ is the stacked vector of all agents’ decision variables, $x^{\star}(p)$ is the generalized Nash equilibrium (GNE) of the game for a given parameter vector $p\in\mathbb{R}^{n_{p}}$, $n_{p}\geq 0$, where by $n_{p}=0$ we mean that no parameter enters (1). In (1), $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is the objective of agent $i$, $g:\mathbb{R}^{n}\to\mathbb{R}^{n_{g}}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{n_{h}}$ define, respectively, (shared) inequality and equality constraints, and $\ell_{i}\in(\mathbb{R}\cup\{-\infty\})^{n_{i}}$, $u_{i}\in(\mathbb{R}\cup\{+\infty\})^{n_{i}}$ define the (local) box constraints for each player $i$. Moreover, $J:\mathbb{R}^{n}\times\mathbb{R}^{n_{p}}\to\mathbb{R}$ is the objective function used to shape the desired equilibrium behavior of the game, and $P\subseteq\mathbb{R}^{n_{p}}$ is a given feasible set for the parameter vector $p$, such as the hyper-box $P=\{p\in\mathbb{R}^{n_{p}}:\ \ell_{p}\leq p\leq u_{p}\}$ where $\ell_{p},u_{p}$ are (possibly infinite) bounds. Problem (1) is also referred to as an optimistic equilibrium of Single-Leader-Multi-Follower (Stackelberg) game [1], with the leader optimizing $p$ and the followers playing the GNEP parameterized by $p$.
A GNE $x^{\star}(p)$ is a vector for which no agent can reduce their cost given the strategies of the other players and feasibility constraints, i.e.,
@@
where, with a slight abuse of notation, we denoted by $\mathop{\rm col}\nolimits(x_{i},x_{-i})$ the vector obtained by collecting the components of $x_{i}$ and $x_{-i}$ in the right order, and $X_{i}(x_{-i}^{\star},p)\subseteq\mathbb{R}^{n_{i}}$ is the feasible set of player $i$ defined by the constraints in (1) given the strategies $x_{-i}^{\star}(p)$ of the other players. We highlight special cases of (1) in the next sections.
@@ 3 KKT Conditions and Nonlinear Least-Squares Formulation
Let us assume that $f_{i}$, $g$, and $h$ are continuously differentiable with respect to $x_{i}$ and that suitable constraint qualifications hold, such as the Linear Independence Constraint Qualification (LICQ) condition at the GNE $x^{\star}(p)$ [^1] The LICQ condition is verified if the gradients of the active constraints at $x^{\star}(p)$ $\bigl\{\nabla_{x_{i}}g_{j}(x^{\star},p):j\in\mathcal{A}(x^{\star})\bigr\}\;\cup\;\bigl\{\nabla_{x_{i}}h_{k}(x^{\star},p)\bigr\}\;\cup\;\bigl\{\pm e_{k}:x_{i,k}^{\star}=\ell_{i,k}\text{ or }u_{i,k}\bigr\}$ are linearly independent, where $e_{k}$ is the $k$-th unit vector of appropriate dimension, $\mathcal{A}(x^{\star})=\{j:g_{j}(x^{\star},p)=0\}$ is the set of active inequality constraints at $x^{\star}$, and $x_{i,k}$, $\ell_{i,k}$, and $u_{i,k}$ are the $k$-th components of $x_{i}$, $\ell_{i}$, and $u_{i}$, respectively.. Then, the necessary KKT conditions of optimality for each player $i$ read as follows [32, Theorem 12.1]:
@@
where $\lambda_{i}\in\mathbb{R}^{m_{i}}$ denotes the vector of Lagrange multipliers of player $i$ associated with the shared inequality constraints $g(x,p)\leq 0$, $\mu_{i}$ the multipliers corresponding to the shared equality constraints $h(x,p)=0$, and $v_{i}$, $y_{i}$ the multipliers associated with bound constraints on $x_{i}$.
@@
where the last three residuals encode the primal feasibility (3c), dual feasibility (3d), and complementary slackness conditions (3e) using the Fischer–Burmeister nonlinear complementarity problem (NCP) function $\phi:\mathbb{R}\to\mathbb{R}$ [18, 15] componentwise
@@
For conditions of existence of a GNE for a given $p$, such as nonemptiness and compactness of the set of feasible decision vectors, convexity assumptions on $f$, $g$, and linearity of $h$, and constraint qualifications, we refer the reader to [17].
@@ 3.2 Variational GNE
Variational GNEs (vGNEs) can be obtained by simply enforcing that the Lagrange multipliers associated with the shared constraints are the same for all players, i.e., by replacing $\{\lambda_{i}\}$ with a single vector $\lambda$ and $\{\mu_{i}\}$ with a single vector $\mu$, which further reduces the number of variables in the zero-finding problem (5).
@@ 3.3 Game design
[…]
where $z=\mathop{\rm col}\nolimits(x,\nu)$. Problem (7) can be solved via standard nonlinear programming methods with respect to both $z$ and $p$.
@@ 3.4 Non-smooth regularization
in the game-design objective $J$; for example, to find a sparse GNE solution $x^{\star}$, by choosing a sufficiently large $\alpha_{1}$. When $\alpha_{1}>0$, due to the non-differentiability of the $\ell_{1}$-norm at zero, we can reformulate the problem by splitting $x$ into positive and negative parts, $x=x_{p}-x_{m}$, with $x_{p},x_{m}\geq 0$ and apply the result in [5, Corollary 1] to minimize
@@
$J_{\rm reg}(x,p)=$ $\alpha_{1}\left(\mathbf{1}^{\top}(x_{p}+x_{m})+\frac{\alpha^{2}}{2}(\|x_{p}\|_{2}^{2}+x_{m}\|_{2}^{2})\right)$
@@
$\min_{p,x_{p},x_{m},\nu}$ $\frac{1}{2}\left\|\begin{bmatrix}\alpha_{3}x_{p}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \alpha_{3}x_{m}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \sqrt{\rho}\ R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu))\end{bmatrix}\right\|_{2}^{2}$ (10)
@@ 4 Linear Quadratic Games
where $x^{\star}$ is the generalized Nash equilibrium of the Linear-Quadratic (LQ) game with $N$ players, $(Q^{i},c^{i},F^{i})$ define the quadratic cost function for agent $i$, with $Q^{i}=(Q^{i})^{\top}\in\mathbb{R}^{n\times n}$, and the diagonal block $Q^{i}_{ii}\succeq 0$, $c^{i}\in\mathbb{R}^{n}$, $F^{i}\in\mathbb{R}^{n\times n_{p}}$, matrices $A\in\mathbb{R}^{n_{g}\times n}$, $b\in\mathbb{R}^{n_{g}}$, $S\in\mathbb{R}^{n_{g}\times n_{p}}$ define the (shared) linear inequality constraints, $A_{\mathrm{eq}}\in\mathbb{R}^{n_{h}\times n}$, $b_{\mathrm{eq}}\in\mathbb{R}^{n_{h}}$, $S_{\mathrm{eq}}\in\mathbb{R}^{n_{h}\times n_{p}}$ define the (shared) linear equality constraints.
@@
or a convex quadratic function
@@
Special cases are: ($i$) $J_{\rm PWA}(x,p)=J_{\rm Q}(x,p)\equiv 0$ for no game design; ($ii$) $J_{\rm PWA}(x,p)=\|x-x_{\rm des}\|_{\infty}$ (i.e., $n_{J}=1$, $n_{1}=2n$, $D_{1}=[I\ -I]^{\top}$, $E_{1}=0$, $h_{1}=\mathop{\rm col}\nolimits(-x_{\rm des},x_{\rm des})$), or $J_{\rm PWA}(x,p)=\|x-x_{\rm des}\|_{1}$ (i.e., $n_{J}=n$, $n_{j}=2$, $D_{j,1}=e_{j}^{\top}$, $D_{j,2}=-e_{j}^{\top}$, $E_{j,1}=E_{j,2}=0$, $h_{j,1}=-x_{\rm des,j}$, $h_{j,2}=x_{\rm des,j}$), or $J_{\rm Q}(x,p)=\frac{1}{2}\|x-x_{\rm des}\|_{2}^{2}$ for inverse game problems; ($iv$) the social welfare quadratic objective $J_{\rm Q}(x,p)=\sum_{i=1}^{N}f_{i}(x,p)$.
@@
where $Q^{i}_{i}$ collects the rows of $Q^{i}$ corresponding to player $i$’s variables, $\bar{A}_{i}$ collects the columns and rows of $\bar{A}$ where $x_{i}$ is involved, $A_{\mathrm{eq},i}$ is defined similarly, and $\lambda_{i}$, $\mu_{i}$, are the Lagrange multipliers associated with the inequality and equality constraints, respectively, where the variables in $x_{i}$ appear.
By stacking the KKT conditions (13) for all players, we can rewrite the generalized Nash equilibrium conditions as a single mixed-integer linear program (MILP) using standard big-M constraints to model the complementarity conditions (see, e.g., [25, Proposition 2]):

### 18 Mar 2026 v3

@@ Abstract
NashOpt is an open-source Python library for computing and designing generalized Nash equilibria (GNEs) in noncooperative games with shared constraints and real-valued decision variables. The library exploits the joint Karush-Kuhn-Tucker (KKT) conditions of all players to handle both general nonlinear GNEs and linear–quadratic games, including their variational versions. Nonlinear games are solved via nonlinear least-squares formulations, relying on JAX for automatic differentiation. Linear-quadratic GNEs are reformulated as mixed-integer linear programs, enabling efficient computation of multiple equilibria. The framework also supports inverse-game and Stackelberg game-design problems. The capabilities of NashOpt are demonstrated through several examples, including noncooperative game-theoretic control problems of linear quadratic regulation and model predictive control. The library is available at https://github.com/bemporad/nashopt.
@@ 1 Introduction
Generalized games describe situations where several agents take decisions by optimizing individual objective functions under local and coupling constraints on their decision variables. Examples include traffic participants using the same roads, wireless systems sharing limited bandwidth, energy systems where prosumers share transmission capacity and balance constraints [11, 40], or firms competing in a market [3]. The solution concept for these problems, in which the cost function and the feasible set of each agent depends on the decisions taken by the other agents, is the generalized Nash equilibrium (GNE), a state in which no agent has an incentive or the possibility to change their decision [2].
GNEPs have been widely studied in the literature, with various algorithms proposed for their solution, including splitting-based operator frameworks [41, 3], approaches built on augmented Lagrangian formulations [36, 28, 9, 29], interior-point optimization strategies [13], Newton-type algorithms [17], distributed algorithms [41, 39, 3], and learning-based methods only relying on measuring best responses [15]. For the specific class of GNEPs in which the shared constraints are linear, efficient solution methods exist for linear objectives [14, 21], multiparametric quadratic programming for strictly convex quadratic objectives [25], and distributed payoff-based algorithms for convex games [38].
@@ 1.1 Contribution
Motivated by the relative scarcity of software packages for solving rather general classes of GNEPs and their application to game-theoretic control problems, in this paper we describe the mathematical formulations of our package NashOpt, a Python library for solving GNEPs based on satisfying the KKT conditions of each player jointly. The library supports nonlinear GNEPs, which are solved via nonlinear least-squares methods, and linear-quadratic GNEPs, which can be solved more efficiently via mixed-integer linear programming (MILP) and can be used to enumerate multiple equilibria for the same game corresponding to different combinations of active constraints at optimality. We also address game-design problems, where a central authority can optimize some parameters of the game to achieve a desired equilibrium in accordance with the optimistic Single-Leader-Multi-Follower (SLMF) Stackelberg game setting [1]. Moreover, we leverage the GNE solution methods to solve game-theoretic control problems, including non-cooperative linear quadratic regulation (LQR) and model predictive control (MPC) problems [24, 31]. The main purpose of the tool is to analyze and predict the effect of different agents pursuing their own objectives in a shared environment, and to design the game to promote desired equilibrium behaviors.
@@
The library is open-source and can be easily installed via pip install nashopt. We refer the reader to the online documentation at https://github.com/bemporad/nashopt for a detailed description of the syntax and usage of the library.
@@ 2 Problem Formulation
where $x_{i}\in\mathbb{R}^{n_{i}}$ collects the decision variables optimized by agent $i$, $i=1,\ldots,N$, $x_{-i}$ collects those of all other agents, $x=\mathop{\rm col}\nolimits(x_{1},\ldots,x_{N})\in\mathbb{R}^{n}$ is the stacked vector of all agents’ decision variables, $x^{\star}$ is constrained to be a generalized Nash equilibrium (GNE) of the game corresponding to the parameter vector $p\in\mathbb{R}^{n_{p}}$, $n_{p}\geq 0$, where by $n_{p}=0$ we mean that no parameter enters (1). In (1), $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is the objective of agent $i$, $g:\mathbb{R}^{n}\to\mathbb{R}^{n_{g}}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{n_{h}}$ define, respectively, shared inequality and equality constraints, and $\ell_{i}\in(\mathbb{R}\cup\{-\infty\})^{n_{i}}$, $u_{i}\in(\mathbb{R}\cup\{+\infty\})^{n_{i}}$ define the (local) box constraints for each player $i$. Moreover, $J:\mathbb{R}^{n}\times\mathbb{R}^{n_{p}}\to\mathbb{R}$ is the objective function used to shape the desired equilibrium of the game, and $P\subseteq\mathbb{R}^{n_{p}}$ is a given feasible set for the parameter vector $p$, such as the hyper-box $P=\{p\in\mathbb{R}^{n_{p}}:\ \ell_{p}\leq p\leq u_{p}\}$ where $\ell_{p},u_{p}$ are (possibly infinite) bounds. Problem (1) is also referred to as an optimistic equilibrium of Single-Leader-Multi-Follower (Stackelberg) game [1], with the leader optimizing $p$ and the followers playing the GNEP determined by $p$.
For any given $p$, a GNE $x^{\star}$ is a vector for which no agent can reduce their cost given the strategies of the other players and feasibility constraints, i.e.,
@@
where, with a slight abuse of notation, we denoted by $\mathop{\rm col}\nolimits(x_{i},x_{-i})$ the vector obtained by collecting the components of $x_{i}$ and $x_{-i}$ in the right order, and $X_{i}(x_{-i}^{\star},p)\subseteq\mathbb{R}^{n_{i}}$ is the feasible set of player $i$ defined by the constraints in (1) given the strategies $x_{-i}$ of the other players. We highlight special cases of (1) in the next sections.
@@ 3 KKT Conditions and Nonlinear Least-Squares Formulation
Let us assume that $f_{i}$, $g$, and $h$ are continuously differentiable with respect to $x_{i}$ and that suitable constraint qualifications hold, such as the Linear Independence Constraint Qualification (LICQ) condition at the GNE $x^{\star}(p)$ [^1] The LICQ condition is verified if the gradients of the active constraints at $x^{\star}(p)$ $\bigl\{\nabla_{x_{i}}g_{j}(x^{\star},p):j\in\mathcal{A}(x^{\star})\bigr\}\;\cup\;\bigl\{\nabla_{x_{i}}h_{k}(x^{\star},p)\bigr\}\;\cup\;\bigl\{\pm e_{k}:x_{i,k}^{\star}=\ell_{i,k}\text{ or }u_{i,k}\bigr\}$ are linearly independent, where $e_{k}$ is the $k$-th unit vector of appropriate dimension, $\mathcal{A}(x^{\star})=\{j:g_{j}(x^{\star},p)=0\}$ is the set of active inequality constraints at $x^{\star}$, and $x_{i,k}$, $\ell_{i,k}$, and $u_{i,k}$ are the $k$-th components of $x_{i}$, $\ell_{i}$, and $u_{i}$, respectively.. Then, the necessary KKT conditions of optimality for each player $i$ read as follows [34, Theorem 12.1]:
@@
where $\lambda_{i}\in\mathbb{R}^{m_{i}}$ denotes the vector of Lagrange multipliers of player $i$ associated with the shared inequality constraints $g(x,p)\leq 0$, $\mu_{i}$ the multipliers corresponding to the shared equality constraints $h(x,p)=0$, and $v_{i}$, $y_{i}$ the multipliers associated with bound constraints on $x_{i}$. Note that the problem setting in (1) could be immediately extended to allow for player-specific constraints with coupling of the form $g_{i}(x,p)\leq 0$, $h_{i}(x,p)=0$ in the KKT system (3).
@@
where the last three residuals encode the primal feasibility (3c), dual feasibility (3d), and complementary slackness conditions (3e) using the Fischer-Burmeister nonlinear complementarity problem (NCP) function $\phi:\mathbb{R}\to\mathbb{R}$ [19, 16] componentwise
@@
For conditions of existence of a GNE for a given $p$, such as nonemptiness and compactness of the set of feasible decision vectors, convexity assumptions on $f$, $g$, and linearity of $h$, and constraint qualifications, we refer the reader to [18]. We remark that only relying on the joint KKT conditions (3) to characterize GNEs would require additional regularity assumptions. In NashOpt, these conditions are not verified numerically by the software. Consequently, when the assumptions above are violated, solving the KKT system should be interpreted as a heuristic computational approach for computing stationary points of the joint KKT system that, however, are not guaranteed a priori to be GNEs. On the other hand, in the case a zero residual $R$ is achieved within a given tolerance at a certain $x^{\star}$, whether this corresponds to a GNE or not can be immediately checked a posteriori by verifying that, for each player $i$, the best responses $\bar{x}_{i}(x^{\star}_{-i},p)\approx x_{i}^{\star}$, or that the associated value function $f_{i}(\mathop{\rm col}\nolimits(\bar{x}_{i},x^{\star}_{-i}),p)\approx f_{i}(x^{\star},p)$.
@@ 3.2 Variational GNE
Variational GNEs (vGNEs) are obtained by simply enforcing that the Lagrange multipliers associated with the shared constraints are the same for all players, i.e., by replacing $\{\lambda_{i}\}$ with a single vector $\lambda$ and $\{\mu_{i}\}$ with a single vector $\mu$, which further reduces the number of variables in the zero-finding problem (5).
@@ 3.3 Game design
$\begin{array}[]{rl}\min_{p,z}\quad&J([I\ 0]z,p)\\ \text{s.t. }&R(z,p)=0{,\ p\in P}.\end{array}$ (7)
Problem (7) can be solved via constrained nonlinear programming methods with respect to both $z$ and $p$.
@@ 3.4 Non-smooth regularization
in the game-design objective $J$; for example, we can promote sparsity of the GNE $x^{\star}$ by choosing a sufficiently large $\alpha_{1}$. When $\alpha_{1}>0$, due to the non-differentiability of the $\ell_{1}$-norm at zero, we can reformulate the problem by splitting $x$ into positive and negative parts, $x=x_{p}-x_{m}$, with $x_{p},x_{m}\geq 0$ and apply the result in [5, Corollary 1] to minimize
@@
$J_{\rm reg}(x,p)=$ ${\alpha_{1}\mathbf{1}^{\top}(x_{p}+x_{m})+\alpha_{2}(\|x_{p}\|_{2}^{2}+\|x_{m}\|_{2}^{2})}$
@@
$\min_{p,x_{p},x_{m},\nu}$ $\frac{1}{2}\left\|\begin{bmatrix}\alpha_{3}x_{p}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \alpha_{3}x_{m}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \sqrt{\rho}\ R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu),p)\end{bmatrix}\right\|_{2}^{2}$ (10)
@@ 4 Linear Quadratic Games
where $x^{\star}$ is the generalized Nash equilibrium of the Linear-Quadratic (LQ) game with $N$ players, $(Q^{i},c^{i},F^{i})$ define the quadratic cost function for agent $i$, with $Q^{i}=(Q^{i})^{\top}\in\mathbb{R}^{n\times n}$, and the diagonal block $Q^{i}_{ii}\succeq 0$, $c^{i}\in\mathbb{R}^{n}$, $F^{i}\in\mathbb{R}^{n\times n_{p}}$, matrices $A\in\mathbb{R}^{n_{g}\times n}$, $b\in\mathbb{R}^{n_{g}}$, $S\in\mathbb{R}^{n_{g}\times n_{p}}$ define the shared linear inequality constraints, $A_{\mathrm{eq}}\in\mathbb{R}^{n_{h}\times n}$, $b_{\mathrm{eq}}\in\mathbb{R}^{n_{h}}$, $S_{\mathrm{eq}}\in\mathbb{R}^{n_{h}\times n_{p}}$ define the shared linear equality constraints.
@@
or the convex quadratic function
@@
Special cases are: ($i$) $J_{\rm PWA}(x,p)=J_{\rm Q}(x,p)\equiv 0$ for no game design; ($ii$) $J_{\rm PWA}(x,p)=\|x-x_{\rm des}\|_{\infty}$ (i.e., $n_{J}=1$, $n_{1}=2n$, $D_{1}=[I\ -I]^{\top}$, $E_{1}=0$, $h_{1}=\mathop{\rm col}\nolimits(-x_{\rm des},x_{\rm des})$), or $J_{\rm PWA}(x,p)=\|x-x_{\rm des}\|_{1}$ (i.e., $n_{J}=n$, $n_{j}=2$, $D_{j,1}=e_{j}^{\top}$, $D_{j,2}=-e_{j}^{\top}$, $E_{j,1}=E_{j,2}=0$, $h_{j,1}=-x_{\rm des,j}$, $h_{j,2}=x_{\rm des,j}$), or $J_{\rm Q}(x,p)=\frac{1}{2}\|x-x_{\rm des}\|_{2}^{2}$ for inverse game problems; (iiiiii) the social welfare quadratic objective $J_{\rm Q}(x,p)=\sum_{i=1}^{N}f_{i}(x,p)$.
@@
where $Q^{i}_{i}$ collects the rows of $Q^{i}$ corresponding to player $i$’s variables, $\bar{A}_{i}$ collects the columns and rows of $\bar{A}$ where $x_{i}$ is involved, $A_{\mathrm{eq},i}$ is defined similarly, and $\lambda_{i}$, $\mu_{i}$, are the Lagrange multipliers associated with the inequality and equality constraints, respectively, where the variables in $x_{i}$ appear. Note that the LQ problem (11) could be extended to handle player-specific linear constraints with coupling of the form $A^{i}x\leq b^{i}+S^{i}p$, $A_{\mathrm{eq}}^{i}x=b_{\mathrm{eq}}^{i}+S_{\mathrm{eq}}^{i}p$, by changing (13) accordingly.
By stacking the KKT conditions (13) for all players, we can rewrite the generalized Nash equilibrium conditions as a single mixed-integer linear program (MILP) using standard big-M constraints to model the complementarity conditions (see, e.g., [26, Proposition 2]):